# 2]Huawei Noah’s Ark Lab

### v1

@@ Abstract
keywords:
lane detection, self-supervised learning, autonomous driving
@@ 1 Introduction
Lane detection is a fundamental task in any autonomous driving system requiring reasoning about the shape and position of marked lanes and is of great importance for path planning, steering of vehicles and line keeping for the autonomous system. It has been extensively studied in computer vision. Given an image of road scene, the objective of lane detection is to estimate the position lane paints with the best possible accuracy. Existing approaches for lane detection focus on developing discriminative feature representations to classify whether each pixel represents the lane and assign it to its respective instance [1, 2], or explicitly learn from pre-defined proposals and perform a detection task [3, 4, 5], both in a supervised fashion.
Nevertheless, there are still challenges to address in lane detection, including issues related to data availability and cross-domain generalization. Real-world driving scenes often undergo dramatic changes due to different camera sensors, road type, changes in illumination and background clutter. For example, a lane detector trained on the data collected in US west coast would have problem to detect the lanes at London Piccadilly circus. This severely limits their scalability. A naive solution might resolve this problem but would require more data on different types of roads, with accurate annotations even in occluded scenarios, inevitably increasing the annotation costs.
@@
In this paper, we study an self-supervised learning strategy, dubbed LaneCorrect, to remedy above issue (see Figure 1). Conventional unsupervised learning methods attempt to detect lanes by using hand-crafted feature and curve fitting (e.g., hough transform [6] and B-spline fitting [7]), but with little success. Inspired by [8], we propose to detect noisy lane paints by leveraging the large-scale LiDAR point cloud in an unsupervised manner. The special material of road markers paint is always designed to reflect enough vehicle lights to be seen even in the poor light condition. It also shows a distinctive reflectivity difference between bare pavement and lane paint in LiDAR point cloud. Taking usage of this reflectivity distortion around the lane, we can extract candidate lane instances from the 3D points by using the DBSCAN [9] alongside RANSAC [10], and then predict the lanes in the 2D image plane by projecting the 3D points. We can use such noisy lane prediction as pseudo labels to train a naïve lane detector. However, the noisy labels may impact the quality of the learned lane detector. Hence, we develop a self-supervised algorithm to leverage geometric symmetries of lanes lines and reduce the perturbations of noisy label.
In view of the geometric consistency behind the low-density separation assumption [11], i.e., data points of the same cluster are likely to be of the same decision boundary, we train a self-supervised lane correction (LaneCorrect) model with the RGB image and its noisy label cues as inputs, producing a corrected label without any knowledge of ground-truth annotation. Since inductive geometric symmetries are inherent characteristics of lane annotation, multiple disturbances of the same lane label can be viewed as the multiple additive Gaussian noises applied to the same noise-free annotation. Specifically, we perturb the original pseudo label with two different augmentations and enable the network to be trained and functioned as a correction with a consistency loss. Both predictions are corrected by the network so one prediction can be utilized as another prediction’s label and vice versa. A reconstruction loss is also added to avoid collapsing to trivial solutions [12]. Furthermore, the lane mask pooling followed by a contrastive loss are augmented in the feature representation for instance similarity learning. With the self-supervised pre-trained model, we distill it to train a student network on arbitrary target lane detection dataset (e.g.TuSimple) without touching its ground-truth label, to predict the self-supervised trained model’s representation of the same image. Note we do not rely on LiDAR at the inference phase and the noisy label input to our LaneCorrect model comes from the prediction of the naïve lane detector.
@@ 2 Related work
Unsupervised lane detection and point cloud segmentation. Early works on lane detection are based on unsupervised methods [13] with handcrafted features. They show poor performance and only tackle simple scenes and obvious lanes [6, 7]. Existing unsupervised point cloud segmentation methods can be categorized into four groups: edge-based [14, 15, 16], region growing [17, 18, 19, 20], model fitting [21, 22] and clustering-based [9, 23, 24, 25, 26, 9]. In this paper, we investigate distinctive intensity of lanes among surrounding environment under LiDAR points and propose an unsupervised 3D lane segmentation approach. We thus obtain the initial lane predictions in 2D plane by projecting the segmented 3D points.
@@
Supervised lane detection Most existing lane detection methods are based on dense prediction approach [27, 2, 28], which treat lane detection as a pixel-level segmentation task. Recently there is a surge of interest in proposal-based methods [29, 3, 4, 5, 30, 31] to perform efficient lane detection. In addition, there are a few row-wise detection [32] and parametric prediction [33] based methods in the literature demonstrate their superiority on lane detection problem.
@@
Learning with noisy annotation. Most existing works on training with noisy annotation employ the strategies of selecting a subset of clean labels [34, 35, 36] or leverage the output predictions of the network to correct the loss [34, 37, 38]. In this work, we introduce a novel self-supervised training scheme that automatically correct the lane labels by learning consistency and instance awareness from the geometry translation and rotation noise.
Leveraging unlabeled data. Recently, some works have explored lane detection algorithms in semi-supervised or unsupervised form. [39] proposed a UDA method from synthetic data to unlabeled target domain. [40] proposed a semi-supervised lane detection method in Hough Transform loss. To utilize native autonomous driving scenes data, we propose a novel method in leveraging unlabeled data with the aid of lidar clues.
@@
Also, contrastive learning has recently shown great promise [41, 42, 12, 43, 44, 45, 46, 47, 48, 49] in self-supervised representation learning. Most of the works focus on the image-level representations. However, there has been a increasing interest in learning the instance similarity which is more effective in downstream tasks such as detection and segmentation [50, 51, 52]. In this paper, we propose an instance similarity learning strategy in the representation level alongside consistency learning to pre-train a lane predictions model without human labels.
@@ 3.2 Unsupervised 3D lane segmentation
We denote the input 3D point clouds as $P=\{(p_{x},p_{y},p_{z},p_{i})\}$, where $p_{x},p_{y}$ and $p_{z}$ represent the 3D point coordinates in LiDAR frame, and $p_{i}$ represents the intensity value of this point. First, we use region growing strategy [17] to segment all the input point cloud $P$, and obtain the ground point cloud $P_{G}\subset P$. It is worth mentioning that lane paintings or road surface markings have distinctive intensity $p_{i}$ among their surrounding environment, i.e., asphalt or cement surface with low intensity for the reason that they are painted with special materials to be clearly witnessed even during the night. Therefore, this property leads to an intensity distortion around the lane in the 3D point space. Based on this prior, we filter the ground point clouds by setting a minimum threshold $\tau$ at the intensity to get the preliminary lane candidate points $P_{L}\subset P_{G}$.
For all lane candidate points, to generate 3D lane instance, we use DBSCAN [9] to cluster the segmented candidate points into $k$ clusters $\{C_{i}\}_{i}^{k}$, where the auto-defined $k$ represents the number of lane instances segmented by our method.
Finally, to reduce the influence of clustering noise on lane fitting, we used RANSAC [10] to perform curve fitting on each cluster $C_{i}$ to obtain 3D lane proposals $Y=\{(p_{x},p_{y},p_{z})\}$. These 3D lane instances $Y$ are projected to 2D frames to generate 2D lane coordinates $y=\{(p_{u},p_{v})\}$, where $p_{u}$ and $p_{v}$ denote the 2D pixel coordinates on 2D image plain.
@@ 3.3 Self-supervised lane correction network
Consistency regularization module. We denote the noisy pseudo annotation and its counterpart, i.e., the potential ground-truth annotation as $\tilde{y}$ and $y_{gt}$ respectively, where $\tilde{y}\in\{0,1\}^{H,W}$ and so as $y_{gt}$. And we assume the noise added to the ground-truth lane annotations, which is introduced during curve fitting and 3D-to-2D projection, as a stochastic transformation matrix $s$ distributed as gaussian noise. Our goal is to learn a function $G$, which can predict the corresponding invertible geometrical transformation matrix $s^{T}$ given the input image $X\in\mathbb{R}^{3\times H\times W}$ and the cue of noisy lane labels:
@@
Following [38, 53], a data augmentation method is used for noisy lane annotations. A pair of lane instances augmented with different geometry noises will be generated by applying rotation and translation operations to the single lane annotation. To create supervision for missed and mislabelled situations, some lane instances will also be randomly removed or injected.
@@
where $G_{\theta}$ is online network updated with gradient, and $G_{\phi}$ is the target network that has the same architecture as the online network. The parameters of target network $G_{\phi}$ are an exponential moving average of the online network $G_{\theta}$.
These two sets of corrected lane predictions refer to the same unique ground truth lane annotations on one single image. As a result, corrected lane predictions of one branch of our SLC model can be used to provide supervision for another branch, which can be written as a consistency loss term:
@@
Given merely consistency regularization loss, the model will converge to the trivial solution (e.g., generates all-zero lane instance). In order to prevent model collapse, we introduce a reconstruction loss $\mathcal{L}_{r}$, which enables original pseudo annotation $y$ also being utilized to supervise the online network:
@@ 3.4 Pseudo-label refinement and other details
As noticed, the SLC network accepts both RGB images and pseudo label clues as input. Thus, although SLC predicts noise-free lane detection results, it is infeasible for network to perform further end-to-end inference on downstream datasets. To allevaite this dilemma, we propose a distillation approach to promote our SLC network, as is illustrated in Figure 3. In our pseudo-label refinement stage, pseudo annotations are generated by the naïve lane detector that has been trained on the source dataset as well as our SLC network. The corrected predictions of the SLC model are used as supervision to train a student lane detector. The distilled student is capable to conduct end-to-end inference.
Moreover, this procedure can be easily migrated to downstream datasets. For specifically, we transfer both the naïve lane detector and SLC onto downstream datasets. The naïve lane detector generates noisy but cheap annotations for the SLC, which SLC takes as input and predicts refined lane labels for further distillation. Finally, the distilled student lane detector is evaluated on the target dataset.
@@ 4.1 Experimental setting
We pre-train our LaneCorrect on large-scale Waymo Open dataset [54]. It contains synchronous LiDAR frames and 2D images. To evaluate our proposed method and make fair comparisons, we distill the SLC network on the target domain to train a student lane detector which can be evaluated directly on the testing set. With the help of our self-supervised pre-trained SLC model, the domain gap is alleviated. Note that we do not introduce LiDAR data in the inference phase and no ImageNet pre-trained model is included in our method. Only the supervised counterparts use backbones pre-trained on ImageNet for comparison. The details can be viewed in Table 1.
We conduct evaluations on four datasets: TuSimple [55], CULane [27], LLAMAS [56] and CurveLanes [4]. TuSimple: TuSimple is a widely used dataset targeted to solve the lane detection problem on highways. It includes 3626 training video clips and 2782 test video clips. Only good weather conditions and daytime data are given. CULane: CULane is a traffic lane detection dataset collected in Beijing and released by Chinese University of Hong Kong. This dataset provides 133,235 frames form a 55 hours of videos, which is then be divided into training, validation and test set by 88880, 9675 and 34680. Specially, the test set provides abundant scene, including one normal and 8 challenging categories. LLAMAS: LLAMAS is one of the largest lane detection datasets with over 100k images in highway scenarios. This dataset is not manually annotated. Instead, the annotations of LLAMAS were generated automatically using high definition map projections. The testing set annotations have not been released, so the evaluation results on the test set will be computed by the publisher of LLAMAS, following the CULane metric. CurveLanes: CurveLanes contains 150K images with human annotated lane labels for difficult scenarios in traffic lane detection. In some scenarios of this dataset, the lane detection task is quite complex and challenging due to curve and fork lanes. The details of the datasets can be viewed in Table 2.
@@
Table 2: Details of lane detection datasets utilized.
Implementation details. To generate pseudo 3D lane segmented points and corresponding 2D coordinates, we firstly perform segmentation algorithm on Waymo Open dataset [54]. We only select synchronized 2D images collected by the camera in the front direction and 3D point clouds from the top LiDAR as our inputs. We choose PointLaneNet [3] as our naïve lane detection network and our supervised baseline, with ResNet101 [57] as backbone in the main results. We additionally run experiments with LaneATT [5] as our baseline in ablation to better prove our framework adaptable to arbitrary lane detectors. During SLC network training, the augmentation method consists of random rotation between $-5^{\circ}$ and $5^{\circ}$ and pixel translation of up to $5\%$ of the width of the input image. We apply an adjusting strategy for reconstruction loss weight to prevent the model overfit to pseudo annotations, which has been detailed described in consistency regularization module. The hyper-paremeter $\lambda_{embed}$ is set to $5$ according to ablation study. All other hyper-parameter settings follow PointLaneNet [3], and our whole architecture is conducted on 8 Nvidia V100 GPU cards.
Details of base lane detector. In our experiments, we adopted anchor-based PointLaneNet [3] as our unsupervised naïve lane detector, and our lane prediction loss was formulated accordingly. Now we give more details about our lane detector.
PointLaneNet [3] can simultaneously perform position prediction and lane classification in a single network. Two $1\times 1$ convolution layers are added on the top of the backbone network, specifying the number of output channels equal to $(n+1)$, where $n$ refers to the number of $x$ coordinate offsets $\{\delta x_{1},\delta x_{2},...,\delta x_{n}\}$ (fixed $y$ partitions) relative to the center point of the grid, and $1$ refers to the starting position $(y)$ of the lane. As for classification, after two $1\times 1$ convolution layers to the end of the feature map, the number of output channels is specified equal to 2, indicating whether the lane passes through the grid. Given ground truth $y_{gt}$, the objective is to optimize the multi-part loss:
@@
$\mathcal{L}_{ij}^{cls}(y_{ij}^{cls},~{}y^{cls}_{ij,~{}gt})$ represents the classification confidence loss at anchor $(i,~{}j)$, which is cross-entropy loss between the prediction results and ground truth. For each anchor with the classfication prediction $y^{cls}$ and $y_{gt}^{cls}$, the confidence loss is written as:

### v2

# 2Huawei Noah’s Ark Lab

@@ Abstract
journal: International Journal of Computer Vision https://doi.org/10.1007/s11263-025-02417-3
@@ 1 Introduction
Lane detection is a fundamental task in any autonomous driving system requiring reasoning about the shape and position of marked lanes and is of great importance for path planning, steering of vehicles and line keeping for the autonomous system. Given an image of the street scene, the objective of lane detection is to estimate the position of the lane paints with the best possible accuracy. Existing approaches for lane detection focus on developing discriminative feature representations to classify whether each pixel represents the lane and assign it to its respective instance (Pan et al., 2018a; Hou et al., 2019), or explicitly learn from pre-defined proposals and perform a detection task (Chen et al., 2019b; Xu et al., 2020; Tabelini et al., 2021b), both in a supervised fashion.
Although the existing supervised methods have been shown to be effective in improving the lane detection performance on specific benchmarks over the past five years, lane detection remains an unsolved problem. This is because that the real-world driving scenes often undergo dramatic changes due to different camera sensors, road type, changes in illumination and background clutter. For example, a lane detector trained on the data collected in US west coast would have problem to detect the lanes at London Piccadilly circus. This severely limits their scalability. Training with more data and annotations from different types of roads even on the occluded scenario in a supervised-learning paradigm might resolve this problem, but it would inevitably increase annotation costs.
@@
In this paper, we study an self-supervised learning strategy, dubbed LaneCorrect, to remedy above issue (see Figure 1). Conventional unsupervised learning methods attempt to detect lanes by using hand-crafted feature and curve fitting (e.g., hough transform Zeng et al. (2019) and B-spline fitting Cao et al. (2019)), but with little success. Inspired by Tian et al. (2021), we propose to detect noisy lane paints by leveraging the large-scale LiDAR point cloud in an unsupervised manner. The special material of road markers paint is always designed to reflect enough vehicle lights to be seen even in the poor light condition. It also shows a distinctive reflectivity difference between bare pavement and lane paint in LiDAR point cloud. Taking usage of this reflectivity distortion around the lane, we can extract candidate lane instances from the 3D points by using the DBSCAN Xu et al. (2018) alongside RANSAC Fischler and Bolles (1981), and then predict the lanes in the 2D image plane by projecting the 3D points. We can use such noisy lane prediction as pseudo labels to train a naïve lane detector. However, the noisy labels may impact the quality of the learned lane detector. Hence, we develop a self-supervised algorithm to leverage geometric symmetries of lanes lines and reduce the perturbations of noisy label.
In view of the geometric consistency behind the low-density separation assumption Ouali et al. (2020), i.e., data points of the same cluster are likely to be of the same decision boundary, we train a self-supervised lane correction (LaneCorrect) model with the RGB image and its noisy label cues as inputs, producing a corrected label without any knowledge of ground-truth annotation. Since inductive geometric symmetries are inherent characteristics of lane annotation, multiple disturbances of the same lane label can be viewed as the multiple additive Gaussian noises applied to the same noise-free annotation. Specifically, we perturb the original pseudo label with two different augmentations and enable the network to be trained and functioned as a correction with a consistency loss. Both predictions are corrected by the network so one prediction can be utilized as another prediction’s label and vice versa. A reconstruction loss is also added to avoid collapsing to trivial solutions He et al. (2020). Furthermore, the lane mask pooling followed by a contrastive loss are augmented in the feature representation for instance similarity learning. With the self-supervised pre-trained model, we distill it to train a student network on arbitrary target lane detection dataset (e.g.TuSimple) without touching its ground-truth label, to predict the self-supervised trained model’s representation of the same image. Note we do not rely on LiDAR at the inference phase and the noisy label input to our LaneCorrect model comes from the prediction of the naïve lane detector.
@@ 2 Related work
Unsupervised lane detection and point cloud segmentation. Early works on lane detection are based on unsupervised methods Zhang et al. (2021) with handcrafted features. They show poor performance and only tackle simple scenes and obvious lanes Zeng et al. (2019); Cao et al. (2019). Existing unsupervised point cloud segmentation methods can be categorized into four gro-  ups: edge-based Nguyen and Le (2013); Bhanu et al. (1986); Jiang et al. (1996), region growing Ning et al. (2009); Tóvári and Pfeifer (2012); Dong et al. (2018); Rabbani et al. (2006), model fitting Hough (1962); Fischler and Bolles (1981) and clustering-based Xu et al. (2018); Shahzad et al. (2012); Zhu and Shahzad (2014); Sampath and Shan (2010); Ferraz et al. (2010). In this paper, we investigate distinctive intensity of lanes among surrounding environment under LiDAR points and propose an unsupervised 3D lane segmentation approach. We thus obtain the initial lane predictions in 2D plane by projecting the segmented 3D points.
@@
Supervised lane detection Most existing lane detection methods are based on dense prediction approach Pan et al. (2018b); Hou et al. (2019); Xu et al. (2022), which treat lane detection as a pixel-level segmentation task. Recently there is a surge of interest in proposal-based methods Li et al. (2019); Chen et al. (2019b); Xu et al. (2020); Tabelini et al. (2021b); Zheng et al. (2022); Liu et al. (2021) to perform efficient lane detection. In addition, there are a few row-wise detection Qin et al. (2020) and parametric prediction Tabelini et al. (2021a) based methods in the literature demonstrate their superiority on lane detection problem.
@@
Learning with noisy annotation. Most existing works on training with noisy annotation employ the strategies of selecting a subset of clean labels Nishi et al. (2021); Malach and Shalev-Shwartz (2017); Jiang et al. (2018) or leverage the output predictions of the network to correct the loss Nishi et al. (2021); Patrini et al. (2017); Chen et al. (2019a). In this work, we introduce a novel self-supervised training scheme that automatically correct the lane labels by learning consistency and instance awareness from the geometry translation and rotation noise.
@@
Leveraging unlabeled data. Recently, several studies explore lane detection algorithms in semi-supervised or unsupervised forms. Garnett et al. (2020) proposes a UDA method to transfer from synthetic data to an unlabeled target domain, while Lin et al. (2021) introduces a semi-supervised lane detection method using Hough Transform loss. To better leverage data from native autonomous driving scenes, we propose a novel method that utilizes unlabeled data with the aid of lidar clues. Additionally, contrastive learning has recently shown great promise Hadsell et al. (2006); Chen et al. (2020); He et al. (2020); Wu et al. (2018); Asano et al. (2020); Caron et al. (2019, 2020); Goyal et al. (2021); He et al. (2022); Gao et al. (2022) in self-supervised representation learning. Most of the works focus on the image-level representations. However, there has been an increasing interest in learning the instance similarity which is more effective in downstream tasks such as detection and segmentation Van Gansbeke et al. (2021); Zhang and Maire (2020); Hénaff et al. (2021). In this paper, we propose an instance similarity learning strategy in the representation level alongside consistency learning to pre-train a lane predictions model without human labels.
@@ 3.2 Unsupervised 3D lane segmentation
We denote the input 3D point clouds as $P=\{(p_{x},p_{y},\\ p_{z},p_{i})\}$, where $p_{x},p_{y}$ and $p_{z}$ represent the 3D point coordinates in LiDAR frame, and $p_{i}$ represents the intensity value of this point. First, we use region growing strategy Ning et al. (2009) to segment all the input point cloud $P$, and obtain the ground point cloud $P_{G}\subset P$. It is worth mentioning that lane paintings or road surface markings have distinctive intensity $p_{i}$ among their surrounding environment, i.e., asphalt or cement surface with low intensity for the reason that they are painted with special materials to be clearly witnessed even during the night. Therefore, this property leads to an intensity distortion around the lane in the 3D point space. Based on this prior, we filter the ground point clouds by setting a minimum threshold $\tau$ at the intensity to get the preliminary lane candidate points $P_{L}\subset P_{G}$.
For all lane candidate points, to generate 3D lane instance, we use DBSCAN Xu et al. (2018) to cluster the segmented candidate points into $k$ clusters $\{C_{i}\}_{i}^{k}$, where the auto-defined $k$ represents the number of lane instances segmented by our method.
Finally, to reduce the influence of clustering noise on lane fitting, we used RANSAC Fischler and Bolles (1981) to perform curve fitting on each cluster $C_{i}$ to obtain 3D lane proposals $Y=\{(p_{x},p_{y},p_{z})\}$. These 3D lane instances $Y$ are projected to 2D frames to generate 2D lane coordinates $y=\{(p_{u},p_{v})\}$, where $p_{u}$ and $p_{v}$ denote the 2D pixel coordinates on 2D image plain.
@@ 3.3 Self-supervised lane correction network
Consistency regularization module. We denote the noisy pseudo annotation and its counterpart, i.e., the potential ground-truth annotation as $\tilde{y}$ and $y_{gt}$ respectively, where $\tilde{y},~{}y_{gt}\in\{0,1\}^{H,W}$. We assume the noise added during the curve fitting and 3D-to-2D projection follows the Gaussian distribution with a stochastic transformation matrix $s$. Our goal is to learn a function $G$, which can predict the corresponding invertible geometrical transformation matrix $s^{T}$ given the input image $X\in\mathbb{R}^{3\times H\times W}$ and the cue of noisy lane labels:
@@
Following Chen et al. (2019a); Zorzi et al. (2020), a data augmentation method is used for noisy lane annotations. A pair of lane instances augmented with different geometry noises will be generated by applying rotation and translation operations to the single lane annotation. To create supervision for missed and mislabelled situations, some lane instances will also be randomly removed or injected.
@@
where $G_{\theta}$ is online network updated with gradient, and $G_{\phi}$ is the target network that has the same architecture as the online network. The parameters of target network $G_{\phi}$ are an exponential moving average of those of the online network $G_{\theta}$.
These two sets of corrected lane predictions refer to the same unique ground truth lane annotations on one single image. As a result, corrected lane predictions of the online branch of SLC can be used to provide supervision for the target branch, which can be written as a consistency loss term:
@@
Given merely consistency regularization loss, the mo-  del will converge to the trivial solution (e.g., generates all-zero lane instance). In order to prevent model collapse, we introduce a reconstruction loss $\mathcal{L}_{r}$, which enables original pseudo annotation $y$ also being utilized to supervise the online network:
@@ 3.4 Pseudo-label refinement and other details
As noticed, the SLC network accepts both RGB images and pseudo label clues as input. Thus, although SLC predicts noise-free lane detection results, it is infeasible for network to perform further end-to-end inference on downstream datasets. To alleviate this dilemma, we propose a distillation approach to promote our SLC network, as is illustrated in Figure 3. In our pseudo-label refinement stage, pseudo annotations are generated by a naïve lane detector that has been trained on the source dataset as well as our SLC network (see Figure 3). The noisy results of unsupervised lane segmentation are replaced by these generated pseudo annotation, thus the 3D LiDAR data is not required for further downstream inference. The corrected predictions of the SLC model are used as supervision to train a student lane detector. The distilled student is capable to conduct end-to-end inference.
Moreover, this procedure can be easily migrated to downstream datasets. For specifically, we transfer both the naïve lane detector and SLC onto downstream data-  sets. The naïve lane detector generates noisy but cheap annotations for the SLC, which SLC takes as input and predicts refined lane labels for further distillation. Finally, the distilled student lane detector is evaluated on the target dataset.
@@ 4.1 Experimental setting
We pre-train our LaneCorrect on large-scale Waymo Open dataset (Sun et al., 2020). It contains synchronous LiDAR frames and 2D images. To evaluate our proposed method and make fair comparisons, we distill the SLC network on the target domain to train a student lane detector which can be evaluated directly on the testing set. With the help of our self-supervised pre-trained SLC model, the domain gap is alleviated. Note that we do not introduce LiDAR data in the inference phase and no ImageNet pre-trained model is included in our method. Only the supervised counterparts use backbones pre-trained on ImageNet for comparison. The details can be viewed in Table 1.
We conduct evaluations on four datasets: TuSimple TuSimple (2019), CULane Pan et al. (2018b), LLAMAS Behrendt and Soussan (2019) and CurveLanes Xu et al. (2020). TuSimple: TuSimple is a widely used dataset targeted to solve the lane detection problem on highways. It includes 3626 training video clips and 2782 test video clips. Only good weather conditions and daytime data are given. CULane: CULane is a traffic lane detection dataset collected in Beijing and released by Chinese University of Hong Kong. This dataset provides 133,235 frames form a 55 hours of videos, which is then be divided into training, validation and test set by 88880, 9675 and 34680. Specially, the test set provides abundant scene, including one normal and 8 challenging categories. LLAMAS: LLAMAS is one of the largest lane detection datasets with over 100,000 images with highway scenarios. The annotations of LLAMAS are automatically generated by the high definition map projections. Its test set evaluation is provided by the official LLAMAS server. CurveLanes: CurveLanes contains 150K images with human annotated lane labels for difficult scenarios in traffic lane detection. In some scenarios of this dataset, the lane detection task is quite complex and challenging due to curve and fork lanes. The details of the datasets can be viewed in Table 2.
Table 2: Details of lane detection datasets utilized.
@@
Implementation details. To generate pseudo 3D lane segmented points and corresponding 2D coordinates, we firstly perform segmentation algorithm on Waymo Open dataset Sun et al. (2020). We only select synchronized 2D images collected by the camera in the front direction and 3D point clouds from the top LiDAR as our inputs. We choose PointLaneNet Chen et al. (2019b) as our naïve lane detection network and our supervised baseline, with ResNet101 He et al. (2016) as backbone in the main results. We additionally run experiments with LaneATT Tabelini et al. (2021b) as our baseline in ablation to better prove our framework adaptable to arbitrary lane detectors. During SLC network training, the augmentation method consists of random rotation between $-5^{\circ}$ and $5^{\circ}$ and pixel translation of up to $5\%$ of the width of the input image. We apply an adjusting strategy for reconstruction loss weight to prevent the model overfit to pseudo annotations, which has been detailed described in consistency regularization module. The hyper-paremeter $\lambda_{embed}$ is set to $5$ according to ablation study. All other hyper-parameter settings follow PointLaneNet Chen et al. (2019b), and our whole architecture is conducted on 8 Nvidia V100 GPU cards.
Details of base lane detector. In our experiments, we adopt the anchor-based counterpart PointLaneNet  Chen et al. (2019b) as the naïve lane detector ($\mathcal{L}_{lane}$ in Eq 4). PointLaneNet Chen et al. (2019b) is able to simultaneously perform position prediction and lane classification in a single network. Two $1\times 1$ convolution layers are added on the top of the backbone network, specifying the number of output channels equal to $(n+1)$, where $n$ refers to the number of $x$ coordinate offsets $\{\delta x_{1},\delta x_{2},...,\delta x_{n}\}$ (fixed $y$ partitions) relative to the center point of the grid, and $1$ refers to the starting position $(y)$ of the lane. As for classification, after two $1\times 1$ convolution layers to the end of the feature map, the number of output channels is specified equal to 2, indicating whether the lane passes through the grid. Given ground truth $y_{gt}$, the objective is to optimize the multi-part loss:
@@
$\mathcal{L}_{ij}^{cls}(y_{ij}^{cls},~{}y^{cls}_{ij,~{}gt})$ represents the classification confidence loss at anchor $(i,~{}j)$, which is cross-entropy loss between the prediction results and ground truth. For each anchor with the classification prediction $y^{cls}$ and $y_{gt}^{cls}$, the confidence loss is written as:
@@
Table 4: Comparison with comparative learning-based counterpart on TuSimple testing set. The first column indicates the proportion of labeled data used for training. The efficiency of our method towards data utilization is more pronounced.
Labels Pseudo labels Methods ACC(%$\uparrow$
100% - ERFNet Lin et al. (2021) 93.71
50% Hough Transform s4GAN Mittal et al. (2019) 88.82
50% Hough Transform ERFNet 93.70
50% 3D lane segmentation Ours 94.04
10% Hough Transform s4GAN 86.25
10% Hough Transform ERFNet 93.05
10% 3D lane segmentation Ours 93.58
0% Hough Transform s4GAN -
0% Hough Transform ERFNet -
0% 3D lane segmentation Ours 91.95